OpenStudy (anonymous):

2(j+k)=m Solve for k

zepdrix (zepdrix):

Hmm so we want to isolate \(\large k\). Let's start by dividing both sides by 2,\[\large \frac{2(j+k)}{2}=\frac{m}{2}\] Hmm what happens to the 2's on the left side?

OpenStudy (anonymous):

I probably would have gone a different way about it because I just manage to be difficult about that. If you Foil out 2 into j and k your result will be. 2j + 2k = m Then you can subtract 2j resulting in 2k = m - 2j after this you can divide out the 2 with the k. k = m/2 - j.